# thanks: Equal contributionthanks: Equal contribution

### v1

@@ Abstract
Drawing inspiration from the motility behaviour of microorganisms, we introduce a highly tunable, robotic system self-actuating into the run-and-tumble (RT)-like motion. It comprises two disk-shaped, centimeter-scale programmable robots individually programmed to perform overdamped active Brownian (AB) motion and connected by a rigid rod. The rod is attached to pivot points located on off-centered, mirror-symmetric points on each robot, allowing for its free rotation at both ends. We show that the collective dynamics of this system execute RT-like motion with characteristic sharp tumble events and exponentially distributed run times, similar to those observed in microorganisms. We further quantify emerging dynamics in terms of tumbling frequency and tune it over a wide range of experimental parameters. We also develop a theoretical model that reproduces our experimental results and elucidates the underlying physical mechanisms governing the rich phase behavior of RT motion. Together, these results highlight the importance of our robotic platform in offering valuable insights into the mechanism of motility characteristics in microorganisms.
@@ I INTRODUCTION
Motility is one of the defining features of active and living organisms across all length scales [1, 2, 3]. One prominent example is run-and-tumble (RT) motion, a behavior commonly observed in swimming microorganisms living in a low Reynolds number environment [4, 5, 6, 7, 8]. As the name suggests, these organisms trace relatively straight paths (”runs”) before abruptly changing to a new, randomly chosen direction (”tumbles”). This type of motion is integral to key biological processes such as chemotaxis, nutrient foraging, escaping unfavorable environments, and colonization [9, 10]. Experimental investigations reveal that the RT motion originates through coordinated dynamics of multiple active units shaped as filamentous appendages, called flagella or cilia, decorating the body of microorganisms. These units exhibit intrinsic activity through their rhythmic beating, propelling the organism forward in a series of synchronous and asynchronous cycles [5, 11, 12, 13, 14, 15, 16, 17, 18]. However, the precise mechanisms by which the collective motion of these active units produce motility remains an open question and a topic of active research.
@@
Most organisms exhibiting RT motion lack cognitive abilities or centralized control systems like the brain. Therefore, simple coupling rules between its active components must govern their dynamics. Since run and tumble states are associated with translational and rotational motion, they likely originate from an interplay between active forces and their corresponding moments respectively. Moreover, since they swim in a fluid medium, attempts have been made to investigate the effects of hydrodynamic coupling between beating flagella on motility  [19, 20, 21, 22, 23, 24]. However, there is growing evidence that such couplings are unlikely to play a significant role, especially in the case of biciliated organisms like Chlamydomonas Reinhardtii [25, 26, 27]. These studies indicate that their motility must arise from simple mechanical coupling between active components and the organism’s body. Previous research conducted on microorganisms strongly supports this notion [28, 26, 29, 30]. Thus, it remains to be seen whether an artificial model system that demonstrates and validates the emergence of RT motion based on these mechanisms can be envisaged.
@@
Figure 1: An analog robotic model system executing run-and-tumble (RT) motion: (a) Photograph of the experimental system featuring two coupled robots. The pivot points, $P_{1}$ and $P_{2}$, allow the connecting rod to rotate freely in the horizontal plane. (b) Each robot is programmed to perform overdamped active Brownian (AB) motion with a constant self-propulsion speed $v_{a}$ = 5 cm s-1. Representative trajectories of the free robots are shown for three different values of the rotational diffusion constant, $D_{r}$ (in rad2 s-1). (c) Schematic diagram of the experimental setup, highlighting key variables incorporated into the theoretical model. (d) A typical RT trajectory of the centroid point $C$, for the coupled robot system with $v_{a}$ = 5 cm s-1, $D_{r}$ = 0.06 rad2 s-1, $\delta$ = 3 cm, and $\alpha=90^{\circ}$. The color bar shows the value of the pair angle parameter $\beta$. Black arrows along the trajectory indicate the direction of motion. (e) Zoomed-in views of typical tumble events under conditions of high (left) and low (right) substrate friction. Low friction results in the wheel skidding, introducing significant inertial time delay and causing tumble events to vanish. We increased $v_{a}$ to 20 cm s-1 for low friction case to better illustrate the effect.
@@
In past, attempts have been made to mimic various dynamical features of microorganisms in synthetic systems [31, 32, 33, 34, 35, 36, 37]. Notwithstanding, an artificial analogue system that accurately mimics RT motion with quick, straight run trajectories and slow, sharp tumbles is lacking. While there are reports of RT-like motion in vibrated granular particles [38] and self-propelling camphor boats [39], these studies often feature long tumble durations and very short runs. This behavior contrasts with microorganisms, where runs typically last approximately an order of magnitude longer than tumbles [4, 5, 6]. This discrepancy likely originates from finite inertial effects in these mesoscale synthetic systems [40, 41, 42, 43], limiting their relevance to microscopic organisms that exclusively operate in a low Reynolds number regime. Therefore, for a mesoscale artificial model system to accurately reproduce RT motion, it must similarly operate in an overdamped regime. To the best of our knowledge, no experimental system currently exists that fulfills these conditions while providing tunable statistical features of the RT motion.
To this end, we introduce an experimental model system supported by a theoretical framework executing highly tunable RT-like dynamics. The setup is rather simple and involves two disk-shaped programmable, wheeled robots physically connected by a rigid rod. The connecting rod is free to rotate within the horizontal plane, with its ends pivoted at specific, mirror-symmetric points on each robot. The setup is conceived to model robots as active components and the rod as the body of an organism. The rod’s placement on each robot is arbitrary, serving as an analogy for biochemical processes that internally regulate filamentary connections [28]. Due to their unique propulsion mechanism based upon rolling-without-slipping wheels, each robot executes an overdamped active Brownian (AB) motion at its center with negligible inertial effects. This AB motion is characterized by a propulsion speed $v_{a}$ and rotational diffusion constant $D_{r}$ which is programmed into each robot. The mechanical linkage transforms the active force exerted by each robot combined with its rotational diffusion into a stochastic torque around vertical axes passing through pivot points. This torque acts in opposite directions for two robots, resulting in the coupled system exhibiting self-actuated motion resembling RT dynamics. We characterize the motion using a single parameter, the tumbling frequency, $\lambda$, and show that it can be systematically varied as a function of locations of linkage points. We also present a theoretical model that relies on Langevin dynamics for connected robots treated as overdamped AB particles. Our model also accounts for torques experienced by robots due to the connecting rod and successfully explains the experimental results, providing a deeper understanding of the origin of RT-like motility in microorganisms.
In this paper, we adopt an integrated experimental and theoretical approach, conducting simulations based on our theoretical model alongside the experiments. We use a consistent color scheme throughout: experimental results are shown in blue, and simulation in red whenever they are presented together. It is essential to clarify that we have kept our theoretical model minimal by ignoring detailed experimental features of the propulsion mechanism of the robots. Therefore, the goal of our theoretical model is not to achieve a quantitative match with experiments but to capture qualitative characteristics and general trends.
II Experimental and theoretical model
@@
Our experimental system consists of two two-wheeled, self-propelled, programmable robots, described in our previous studies (Fig. 1(a), also Supplemental Material (SM) section IA) [44, 45]. Each robot is 7.5 cm in diameter and individually programmed to execute in-plane overdamped active Brownian (AB) motion along axes passing through their centers, aligned parallel to their wheels [44]. Both robots are equipped with 3D-printed caps featuring black, triangle-shaped markers, enabling precise tracking of their in-plane positions and orientations. The AB trajectories of the single, non-connected robot are shown in Fig. 1(b) for $v_{a}=5$ cm s-1 and three typical values of $D_{r}$ (see SM movie 1). Central to their overdamped property is the fact that they propel forward using their wheels, which roll without slipping on a flat glass surface covered with a white sheet of paper. This provides sufficient traction for the wheels, making the effects of inertia negligible. Next, we physically connected these robots with a rigid rod of length $\ell=14.5$ cm (Fig. 1(a)). The rod is free to rotate about pivot points ($P_{1}$ and $P_{2}$) located on the left and right sides of the two robots, respectively. Here, we vary $\delta$ and $\alpha$ between $1$ and $3$ cm and $30^{\circ}$ and $150^{\circ}$ in intervals of 1 cm and $30^{\circ}$ respectively. A combination of self-propulsion and constraint force due to the rigid rod generates net torque around vertical axes passing through $P_{1}$ and $P_{2}$. These torques act in opposite handedness for the two robots indicated by curved arrows in Fig. 1(a).
Next, we develop a simplified model details of which are presented in the SM section IIA and IIB. Here, all dissipative forces, such as frictional and resistive, acting on a single robot with translational velocity $\mathbf{v}$ and rotational velocity $\bm{\omega}$ are lumped into generalized expressions:
$\mathbf{F}_{\text{d}}$ $=-\bm{\Gamma}\cdot\mathbf{v},$ (1)
$\bm{\tau}_{\text{d}}$ $=-\Gamma_{\tau}\bm{\omega},$ (2)
where $\bm{\Gamma}$ and $\Gamma_{\tau}$ are the dissipation tensor and dissipation coefficient for translational and rotational motion, respectively. We assume that even when connected, robots continue to operate in an overdamped regime in experiments. We also ignore any dependence of $\bm{\Gamma}$ and $\Gamma_{\tau}$ on $\mathbf{v}$ and $\bm{\omega}$ in the leading order. In experiments, the robot encounters significantly greater frictional forces perpendicular to its polar axis (due to wheel sliding) compared to those in the parallel direction (from rolling wheels). As a result, $\bm{\Gamma}$ assumes a non-scalar matrix form that depends on the robot’s orientation $\mathbf{n}$. The dissipative forces along and perpendicular to $\mathbf{n}$ are expressed as $-\Gamma_{\parallel}\mathbf{v}_{\parallel}$ and $-\Gamma_{\perp}\mathbf{v}_{\perp}$, respectively, where $\mathbf{v}_{\parallel}$ and $\mathbf{v}_{\perp}$ denote the projections of $\mathbf{v}$ parallel and perpendicular to $\mathbf{n}$, and $\Gamma_{\parallel}$ and $\Gamma_{\perp}$ are positive constants.
Based on the above assumptions, our minimal model of the coupled AB robots, schematically sketched in Fig. 1(c), is as follows. Let $\mathbf{n}_{i}$ ($i$ = 1 &\&& 2 for two robots) represent the direction of the active velocity of the $i$th robot, and $\mathbf{s}_{i}$ denote the direction from the pivot point $P_{i}$ to robot center $\mathbf{r}_{i}$. The equations of motion for the robot centers $\mathbf{r}_{i}$ and orientation angles $\phi_{i}$ of $\mathbf{s}_{i}$ read
$\frac{d\mathbf{r}_{i}}{dt}$ $=\mathbf{n}_{i}v_{a}+p_{i}F\left[\gamma\mathbf{n}_{i}(\mathbf{n}_{i}\cdot\mathbf{N}_{\perp})+\frac{1}{\Gamma_{\perp}}\mathbf{N}_{\perp}\right],$ (3a)
$\frac{d\phi_{i}}{dt}$ $=-p_{i}F\delta\frac{1}{\Gamma_{\tau}}\mathbf{s}_{i}\cdot\mathbf{N}+\sqrt{2D_{r}}\eta_{i}(t),$ (3b)
where $\eta_{i}(t)$ is delta-correlated white Gaussian noise for the $i$th robot, with $\eta_{1}$ and $\eta_{2}$ being uncorrelated, and $\mathbf{N}$ is the unit vector normal to the rod and $\mathbf{N}_{\perp}$ is the unit vector along the rod (Fig.  1c). Here $p_{1}=1$, $p_{2}=-1$ and $\gamma=(\Gamma_{\perp}-\Gamma_{\parallel})/(\Gamma_{\perp}\Gamma_{\parallel})$. The force $F$, calculated using the constraint that the distance between the $P_{1}$ and $P_{2}$ remains fixed to the rod length $\ell$, is given by
$F$ $=$ $\mathcal{G}(\theta_{1},\theta_{2})\big{\{}-v_{a}L_{\perp}\big{.}$ (4)
.$\big{.}+\sqrt{2D_{r}}\delta\left[-\cos(\theta_{1}+\alpha)\eta_{1}+\cos(\theta_{2}-\alpha)\eta_{2}\right]\big{\}},$
where $\theta_{i}$ is the orientation angle of $\mathbf{n}_{i}$ with respect to $\mathbf{N}$ and it has the following relation with $\phi_{i}$:
$\phi_{1}=\theta_{1}+\theta+\alpha-\pi,\phi_{2}=\theta_{2}+\theta+\pi-\alpha.$ (5)
The function $\mathcal{G}$ has the following form
$\mathcal{G}(\theta_{1},\theta_{2})=\frac{1}{\gamma K_{\perp}+\frac{\delta^{2}}{\Gamma_{\tau}}M_{\perp}+\frac{2}{\Gamma_{\perp}}}.$ (6)
The equation of motion for the orientation angle $\theta$ of $\mathbf{N}$, details of which are provided in SM section IIB, is given by
$\frac{d\theta}{dt}$ $=$ $\frac{v_{a}}{\ell}L_{\parallel}+\frac{1}{\ell}\left(\gamma K_{\parallel}+\frac{\delta^{2}}{\Gamma_{\tau}}M_{\parallel}\right)F$ (7)
$+\frac{\delta}{\ell}\sqrt{2D_{r}}\left[-\sin(\theta_{1}+\alpha)\eta_{1}+\sin(\theta_{2}-\alpha)\eta_{2}\right].$
where $K_{\perp}$, $M_{\perp}$, $L_{\perp}$, $K_{\parallel}$, $M_{\parallel}$, and $L_{\parallel}$ are functions of $\theta_{1}$ and $\theta_{2}$. Their explicit forms are provided in Equations (32)-(37) in the SM section IIB.
By solving Eq. (7) for $\theta(t)$, we determine $\mathbf{N}$, which is then used in Eqs. (3) to compute the trajectories of $\mathbf{r}_{i}$ and $\phi_{i}$. Subsequently, we obtain the trajectory of the centroid of two robots $\mathbf{r}=(\mathbf{r}_{1}+\mathbf{r}_{2})/2$ indicated by point $C$. We set all parameters to the values used in the experiments, except for the following phenomenological ratios, which are adjusted to best match the experimental results: $\Gamma_{\parallel}/\Gamma_{\perp}=0.1$ and $14\Gamma_{\tau}/(\Gamma_{\parallel}d^{2})=$ 1, where $d=7.5$ cm is the diameter of our robots.
Figure 2: Characterization of RT motion in experiment and simulation: (a, b) Probability distribution of $\beta$ shows a pronounced peak near $\beta$ = 0. Inset: small $\beta$ values correspond to faster dynamics whereas higher values correspond to slow speeds. Black dashed lines at $\beta$ = 60 set an arbitrary threshold to differentiate between run and tumble events. (c, d) The parameter $\beta$ plotted as a function of time $t$. $\tau_{t}$ and $\tau_{\text{run}}$ represent tumble duration and run-time respectively. (e, f) In both experiment and simulation, we observe exponentially decaying run times and unimodal tumble durations. Experimental and simulation parameters are $v_{a}$ = 5 cm s-1, $D_{r}$ = 0.06 rad2 s-1, $\delta$ = 3 cm, and $\alpha=90^{\circ}$.
III RESULTS
@@
We start with parameters $\alpha=90^{\circ}$ and $\delta=3$ cm. When connected robots are set in motion, their centroid, $C$, exhibits dynamics resembling fast, nearly straight trajectories (runs) with occurrences of abrupt sharp turns and sudden halts (tumbles). A typical trajectory is shown in Fig. 1(d) and SM movie 2 for $D_{r}=0.06$ rad2 s-1 and 1.33 rad2 s-1 in SM movie 3. We reproduce these dynamics in simulations by numerically solving Eqs. (3) and (7) of our theoretical model (SM movies 2 and 3). As a vital point, we find it essential to minimize inertial time scales in capturing sudden, sharp tumbling events in experiments. This implies minimizing factors of $m/\Gamma_{\parallel}$, $m/\Gamma_{\perp}$ and $I/\Gamma_{\tau}$ where $m$ and $I$ are the mass and the moment of inertia of each robot respectively (Eqs. 4 and 5 in SM section IIA). To achieve this limit, we performed an experiment on a slippery glass surface coated with coconut oil. Interestingly, we find that a majority of tumbling events become significantly smoother, making them harder to distinguish from run states (See Fig. 1(e) and SM movie 4). In contrast, such smooth turns are rare on a frictional substrate. These findings are in accordance with previous studies in some active matter experiments where inertial effects are known to introduce time delays preventing abrupt changes in direction [40, 41, 42]. We further demonstrate that the RT motion does not occur when individual robots are programmed to perform Brownian dynamics (SM movie 5 and SM section ID), confirming its inherently active nature [46]. Henceforth, all experimental results are obtained from experiments conducted on a frictional surface.
@@
To carry out statistical analysis of the resulting RT motion, we record long-time trajectories comprising hundreds of running and tumbling events for experiments and simulations. In experiments, we eliminate trajectories that experienced disruptions due to boundary walls (see SM section IC for more details). We find that two key parameters effectively describe the RT motion: the pair angle $\beta(t)$ representing the instantaneous difference in the in-plane orientations of two robots (see schematic Fig. 1(c)) and the instantaneous speed $V(t)$ of the centroid $C$. A typical distribution of $\beta$ is shown in Figs. 2(a) and 2(b) for experiment and simulation, respectively. We also find that these parameters are inversely related to each other (inset of Figs. 2(a) and 2(b)) with a cluster of points at high-speed values corresponding to low values $\beta$, which we identify as run states. To differentiate runs from tumbles in a quantitative sense, we set an arbitrary threshold of $\beta=60^{\circ}$, corresponding approximately to the half-maximum of the run peak. The typical behaviour of $\beta$ as a function of time is shown in Figs. 2(c) and 2(d) for experiment and simulation, respectively. By using $\beta=60^{\circ}$ as a reference line, we quantify both the tumbling duration ($\tau_{t}$) and run-time ($\tau_{\text{run}}$) from this time series for both experiment and simulation. The resulting distributions $\tau_{\text{run}}$ show exponential decay in each case (Figs. 2(e) and 2(f)). Encouragingly, such exponential decays are ubiquitous in many microorganisms like swimming bacteria [4], Chlamydomonas Reinhartdii algae [5], and amoeba [6]. Our experiments also reveal that tumbles are sudden events, with $\tau_{t}$ uniformly distributed around its mean value, which is approximately five times smaller than the average $\tau_{\text{run}}$ in experiments, consistent with observations made in real biological systems [4, 5].
@@
Figure 3: Quantifying RT motion in terms of tumbling frequency in experiment and simulation: (a, b) The tumbling frequency, $\lambda$ increases with higher $D_{r}$ and $\delta$. (c, d) The phase diagram of $\lambda$ in $\alpha-D_{r}$ plane for $\delta$ = 3 cm shows that run events become significantly more likely for $\alpha>90^{\circ}$. The region enclosed by the dashed line corresponds to $\lambda\approx 0$. (e, f) $\lambda$ exhibits a critical $D_{r}$ indicated by a black arrowhead beyond which tumbling emerges in the system. (g, h) The run speed, $V_{\text{run}}\approx v_{a}$ for $\alpha\leq 90^{\circ}$ but decreases monotonically for $>90^{\circ}$. Error bars represent standard deviation.

### v2

@@ Abstract
Run-and-tumble (RT) motion is commonly observed in flagellated microswimmers, arising from synchronous and asynchronous flagellar beating. In addition to hydrodynamic interactions, mechanical coupling has recently been recognized to play a key role in flagellar synchronization. To explore this, we design a macroscopic model system that comprises dry, self-propelled robots linked by a rigid rod to model a biflagellated microorganism. To mimic a low Reynolds number environment, we program each robot to undergo overdamped active Brownian (AB) motion. We find that such a system exhibits RT-like behavior, characterized by sharp tumbles and exponentially distributed run times, consistent with real microswimmers. We quantify tumbling frequency and demonstrate its tunability across experimental parameters. Additionally, we provide a theoretical model that reproduces our results, elucidating physical mechanisms governing RT dynamics.
@@
Motility is one of the defining features of active and living organisms across length scales [1, 2, 3]. One prominent example is run-and-tumble (RT) motion commonly seen in swimming microorganisms living in a low Reynolds number environment [4, 5, 6, 7, 8]. Here, organisms trace relatively straight paths (runs) before abruptly changing to a new, randomly chosen direction (tumbles). It is known to originate from coordinated dynamics of multiple active units shaped as filamentous appendages, called flagella or cilia, decorating the body of microorganisms. These units exhibit intrinsic activity through their rhythmic beating, propelling the organism forward in a series of synchronous and asynchronous cycles [5, 9, 10, 11, 12, 13, 14, 15, 16].
@@
Most organisms showing RT motion lack cognitive abilities or centralized control systems like the brain. Therefore, simple coupling rules between its active components must govern their dynamics. Since they swim in a fluid medium, attempts have been made to investigate the effects of hydrodynamic coupling between beating flagella on motility  [17, 18, 19, 20, 21, 22]. However, there is growing evidence that mechanical couplings between active components inside the organism’s body also play an important role. This is especially true for biflagellated organisms like Chlamydomonas reinhardtii [16, 23, 24, 25, 23, 26, 27]. Thus, it remains to be seen whether an artificial model system that demonstrates and validates the emergence of RT motion based on these mechanisms can be envisaged.
@@
Figure 1: (a) Photograph of the experimental system featuring coupled robots. The pivot points, $P_{1}$ and $P_{2}$, allow the connecting rod to rotate freely in the horizontal plane. (b) Typical trajectories of free robots executing overdamped AB motion with $v_{a}$ = 5 cm s-1 for three different values of the rotational diffusion constant, $D_{r}$ (in rad2 s-1). (c) A schematic diagram highlighting key variables incorporated into the theoretical model. (d) A typical RT trajectory of the centroid point $C$ in the direction of black arrows, with $v_{a}$ = 5 cm s-1, $D_{r}$ = 0.06 rad2 s-1, $\delta$ = 3 cm, and $\alpha=90^{\circ}$. The color bar shows the value of $\beta$. (e) Zoomed-in views of typical tumble events under conditions of high (top) and low (bottom) substrate friction. We increased $v_{a}$ to 20 cm s-1 for low friction case to better illustrate the smoothening of the tumble event.
@@
In the past, attempts have been made to mimic various dynamical features of microorganisms in synthetic systems [28, 29, 30, 31, 32, 33, 34]. Notwithstanding, an artificial analog system that accurately mimics RT motion with quick, straight run trajectories and slow, sharp tumbles is lacking. While there are reports of RT-like motion in vibrated granular particles [35] and self-propelling camphor boats [36], these studies often feature long tumble durations and very short runs. This behavior contrasts with microorganisms, where runs are much longer than tumbles [4, 5, 6]. This discrepancy likely originates due to finite inertial effects in these macroscale synthetic systems [37, 38, 39, 40], limiting their relevance to microscopic organisms that exclusively operate in inertialess limit. Therefore, for a macroscale artificial model system to accurately reproduce RT motion, it must similarly operate in an overdamped regime. To the best of our knowledge, no experimental system currently exists that fulfills these conditions while providing tunable statistical features of the RT motion. In this study, we introduce an experimental model system supported by a theoretical framework where a highly tunable RT-like motion spontaneously emerges in mechanically coupled self-propelled robots.
@@
Our experimental system consists of two self-propelled robots connected with a rigid rod as shown in Fig. 1(a). The setup is conceived to model robots as active components and the rod as the body of an organism. Each robot is 7.5 cm in diameter and individually programmed to execute in-plane overdamped active Brownian (AB) motion along axes passing through their centers, aligned parallel to their wheels [41, 42] [also Supplemental Material (SM) Section IA]. This AB motion is characterized by a constant self-propulsion speed $v_{a}$ and the rotational diffusion constant $D_{r}$. Both robots are covered with 3D-printed caps featuring black, triangle-shaped markers, for tracking their in-plane positions and orientations. The AB trajectories of the single, non-connected robot are shown in Fig. 1(b) for three typical values of $D_{r}$ while keeping $v_{a}$ = 5 cm s-1 (see SM movie 1). Central to their overdamped feature is the fact that they propel forward using their wheels, which roll without slipping on a flat glass surface covered with a white sheet of paper. This provides sufficient traction for the wheels, making the effects of inertia negligible. The connecting rod of length $\ell=14.5$ cm is free to rotate about pivot points ($P_{1}$ and $P_{2}$) located on the off-centered, mirror-symmetric points on each robot [Figs. 1(a) &\&& (c)]. Here, $\delta$ is the distance of these points from robots’ centres, and $\alpha$ denotes the angle between the pivot-to-center line and the polarity axis. We vary $D_{r}$, $\delta$ and $\alpha$ between 0.06 - 1.33 rad2s-1, $1$ $3$ cm and $30^{\circ}$ $150^{\circ}$ respectively. $v_{a}$ is kept fixed at 5 cm s-1. A combination of self-propulsion and constraint force due to the rigid rod generates net torque around vertical axes passing through $P_{1}$ and $P_{2}$. These torques act stochastically in opposite handedness for the two robots indicated by curved arrows in Fig. 1(a). To summarize, this setup incorporates the fact that the run and tumble states are associated with translational and rotational motion which originate from an interplay between active forces and their corresponding moments, respectively. In what follows, we show that this simple-minded design is capable of showing RT-like motion.
We now develop a theoretical model for our system (for details see SM Secs. IIA and IIB). In our model, all dissipative forces, including frictional and resistive effects on a single robot with translational and rotational velocity $\mathbf{v}$\&& $\bm{\omega}$, are represented by a dissipative force and torque $\mathbf{F}_{\text{d}}=-\bm{\Gamma}\cdot\mathbf{v}$ and $\bm{\tau}_{\text{d}}=-\Gamma_{\tau}\bm{\omega}$, respectively. Here, $\bm{\Gamma}$ and $\Gamma_{\tau}$ denote the translational dissipation tensor and rotational dissipation coefficient, respectively. We assume that even when connected, robots remain overdamped. We also neglect the leading-order dependence of $\bm{\Gamma}$ and $\Gamma_{\tau}$ on $\mathbf{v}$ and $\bm{\omega}$. In experiments, the friction perpendicular to a robot’s polar axis (from wheel sliding) is significantly greater than the parallel friction (from rolling). Consequently, $\bm{\Gamma}$ adopts a non-scalar form dependent on the robot’s orientation $\mathbf{n}$, with dissipative forces given by $-\Gamma_{\parallel}\mathbf{v}_{\parallel}$ and $-\Gamma_{\perp}\mathbf{v}_{\perp}$, where $\mathbf{v}_{\parallel}$ and $\mathbf{v}_{\perp}$ are the velocity components along and perpendicular to $\mathbf{n}$, and $\Gamma_{\parallel},\Gamma_{\perp}>0$. Using these assumptions, we derive the equations of motion for the orientation $\theta$ of the unit vector $\mathbf{N}$, which is normal to the rod, the orientation angles $\theta_{i}$ of the robots ($i=1,2$) relative to $\mathbf{N}$, and the centroid, $C$, of the system, $\mathbf{r}=(\mathbf{r}_{1}+\mathbf{r}_{2})/2$, where $\mathbf{r}_{i}$ represents the center’s position of $i$th robot [see Fig. 1(c) and Eqs. (30–32) of SM]. All parameters match experimental values, except for two phenomenological ratios adjusted for optimal agreement: $\Gamma_{\parallel}/\Gamma_{\perp}=0.1$ and $14\Gamma_{\tau}/(\Gamma_{\parallel}d^{2})=$ 1, with robot diameter $d=7.5$ cm.
@@
We start with parameters $\alpha=90^{\circ}$ and $\delta=3$ cm. When connected robots are set in motion, the point $C$ exhibits dynamics resembling fast, nearly straight trajectories (runs) with occurrences of abrupt sharp turns and sudden halts (tumbles). A typical trajectory is shown in Fig. 1(d) and SM movie 2 for $D_{r}=0.06$ rad2 s-1 and 1.33 rad2 s-1 in SM movie 3. We reproduce these dynamics in simulations by solving equations of $\theta_{i}$, $\theta$ and $\mathbf{r}_{i}$ of our theoretical model (SM movies 2 and 3). We highlight that it essential to minimize inertial time scales in capturing sudden, sharp tumbling events in experiments. This implies minimizing factors of $m/\Gamma_{\parallel}$, $m/\Gamma_{\perp}$ and $I/\Gamma_{\tau}$, where $m$ and $I$ are the mass and the moment of inertia of each robot respectively (Eqs. 4 and 5 in SM section IIA). To achieve the limit where these factors are non-negligible, we performed an experiment on a slippery glass surface coated with coconut oil. Interestingly, we find that a majority of tumbling events become significantly smoother (See Fig. 1(e) and SM movie 4), making them harder to distinguish from run states. In contrast, such smooth turns are rare on a frictional substrate. These findings are in accordance with previous studies in some active matter experiments where inertial effects are known to introduce time delays preventing abrupt changes in direction [37, 38, 39]. We further demonstrate that the RT motion does not occur when individual robots are programmed to perform Brownian dynamics (SM movie 5 and SM section IB), confirming its inherently active nature [43]. Henceforth, all results are obtained from experiments conducted on a frictional surface.
Figure 2: (a, b) Probability distribution of $\beta$ shows a pronounced peak near $\beta$ = 0. Insets: $\beta$ and $V$ are inversely related to each other for both experiment and simulation. Black dashed lines at $\beta$ = 60 sets a threshold to differentiate between run and tumble events. (c, d) Typical behaviour of $\beta$ as a function of time. $\tau_{t}$ and $\tau_{\text{run}}$ represent tumble duration and run-time respectively. (e, f) In both experiment and simulation, we observe exponentially decaying run times (dashed lines as a guide) and unimodal tumble durations exponentially. Experimental and simulation parameters are $v_{a}$ = 5 cm s-1, $D_{r}$ = 0.06 rad2 s-1, $\delta$ = 3 cm, and $\alpha=90^{\circ}$.
@@
To carry out statistical analysis of the RT motion, we record long-time trajectories comprising hundreds of running and tumbling events for experiments and simulations. In experiments, we eliminate trajectories that experience disruptions due to boundary walls (see SM section IC for more details). We find that two key parameters effectively describe the RT motion: the pair angle $\beta(t)$ representing the instantaneous difference in the in-plane orientations of two robots [see schematic Fig. 1(c)] and the instantaneous speed $V(t)$ of the centroid $C$. A typical distribution of $\beta$ is shown in Figs. 2(a) and 2(b) for experiment and simulation, respectively. We also find that these parameters are inversely related to each other [insets of Figs. 2(a) and 2(b)] with a cluster of points at high $V$ and low $\beta$, which we identify as run states and vice versa for tumbles. To differentiate runs from tumbles quantitatively, we set an arbitrary threshold of $\beta=60^{\circ}$, corresponding to the half-maximum of the run peak. The typical behaviour of $\beta$ as a function of time is shown in Figs. 2(c) and 2(d) for experiment and simulation, respectively. By using $\beta=60^{\circ}$ as a reference line, we quantify tumbling duration ($\tau_{t}$) and run-time ($\tau_{\text{run}}$) from this time series for both experiment and simulation. The resulting distributions $\tau_{\text{run}}$ show exponential decay in each case [Figs. 2(e) and 2(f)]. Encouragingly, such exponential decays are ubiquitous in many microorganisms like swimming bacteria [4], Chlamydomonas Reinhartdii algae [5], and amoeba [6]. Our experiments also reveal that runs last much longer than tumbles [$<\tau_{\text{run}}>=7$ s, $<\tau_{t}>=3.5$ s for data presented in Fig. 2(c)] consistent with observations made in real biological systems [4, 5].
@@
Figure 3: (a, b) The tumbling frequency, $\lambda$ increases with $D_{r}$ and $\delta$. (c, d) The phase diagram of $\lambda$ in $\alpha-D_{r}$ plane for $\delta$ = 3 cm. The region enclosed by the dashed line corresponds to $\lambda\approx 0$. (e, f) $\lambda$ exhibits a critical $D_{r}$ indicated by a black arrowhead beyond which tumbling emerges in the system. (g, h) The run speed, $V_{\text{run}}\approx v_{a}$ for $\alpha\leq 90^{\circ}$ but decreases monotonically for $>90^{\circ}$. Here $D_{r}$ = 0.06 rad2 s-1 and error bars represent standard deviation.